OpenStudy (anonymous):

Can someone help me find the inverse of -20x-16?

OpenStudy (ash2326):

y=-20x-16 express y in terms of x y+16=-20x x=-(y+16)/20

OpenStudy (ash2326):

it can by y or f(x)

OpenStudy (anonymous):

y=-20x-16 x=-20y-16 solve for y 20y=-x-16 y=-(x+16)/20

OpenStudy (anonymous):

do u have a project?

OpenStudy (anonymous):

yea